OpenStudy (anonymous):

Use the slope formula to find the missing coordinate. A( a0 , 0), B( 5, 10) m = 2

OpenStudy (amistre64):

and the slope formula is?

OpenStudy (amistre64):

correct, so what is the issue we are having exactly?

OpenStudy (anonymous):

a0= 3 but it showed incorrect

OpenStudy (anonymous):

A( ? , 0), B( 5, 10) m = 2

OpenStudy (amistre64):

i never actually use the formula itself since i always tend to put the wrong number in the wrong spots :/

OpenStudy (anonymous):

?,0 5,10 x1,y1 x2,y2

OpenStudy (amistre64):

A( a0 , 0), B( 5, 10) m = 2 subtract one point from the other, adn then stack y/x A( a0 , 0) - B( 5, 10) m = 2 ----------- a0-5,-10 -10/(a0-5) = 2 -5 = a0-5 i think a0 = 0

OpenStudy (anonymous):

ok thanks i got one more i need help with

OpenStudy (anonymous):

C(5, -2), D(3, ?) m = 0

OpenStudy (amistre64):

m=0 represent a horizontal line such that all the y values of the points are exactly the same

OpenStudy (anonymous):

so ?=0

OpenStudy (amistre64):

x , y C(5, -2) D(3 , ? ) all the y values for m=0 are the same, since C has y=-2, the D has y= ?

OpenStudy (anonymous):

-2 ???????

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you so much how can i give you a medal or something

OpenStudy (amistre64):

there should be a blue .... thats the one lol

OpenStudy (anonymous):

:)

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

are you still there?

OpenStudy (anonymous):

Lines k, l, m, and n have slopes of , - 4, -1and respectively. Which pairs of lines are perpendicular? k =? l =? What is the formula